OpenStudy (meilendurcer):

Please I really need help with this... Award will be given. 2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 over 8 = 7x−1over 6 24 over 1(3x+6 over 8) = 24 over 1(7x−1 over 6) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (meilendurcer):

\[\frac{ 3x+6 }{ 8 } = \frac{ 7x-1 }{ 6 }\] \[\frac{ 24 }{ 1 } (\frac{ 3x+6 }{ 8 } ) = \frac{ 24 }{ 1 } (\frac{ 7x-1 }{ 6 } )\]

OpenStudy (anonymous):

first of all, this is the biology section, try entering mathematics section next time, your question will be answered faster. secondly, you can go to this site for a second help http://www.freemathhelp.com/factoring-calculator.php

OpenStudy (meilendurcer):

ah thanks.... didn't notice

OpenStudy (anonymous):

it looks like he was correct up to the point where he distributed 4(7x -1) and got 28x - 1. The error was failing to distribute the 4 across BOTH terms of the binomial. Should have been 28x - 4 . You can check your work by substituting the value obtained for the variable back into the original equation. \[\frac{ 3(1)+6 }{ 8} = \frac{ 7(1)-1 }{ 6}\] \[\frac{ 9 }{ 8 } \not equal \frac{ 6 }{ 6 }\] Therefore x = 1 is not the correct solution.

OpenStudy (anonymous):

so heres the right solution 3(3x+6) = 4(7x−1) 9x+18 = 28x−4 -9x -9x 18 = 19x - 4 +4 +4 22 = 19x --------- 19 19 1.15 = x

OpenStudy (anonymous):

@meilendurcer